OpenStudy (anonymous):

Simplify. (5√2 - 1)^2

Nnesha (nnesha):

\[(5\sqrt{2}-1)^2\]is it your question ? or \[(5\sqrt{2-1})^2\]

OpenStudy (anonymous):

yup

Nnesha (nnesha):

which one ?

OpenStudy (anonymous):

the second one

OpenStudy (anonymous):

no the first one sorry

Nnesha (nnesha):

okay \[(5\sqrt{2}-1)^2\] is same as \[(5 \sqrt{2} -1)(5 \sqrt{2} -1)\] do you know foil method ?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

can you teach me?

Nnesha (nnesha):

sure! foil method when you multiply everything which is in first parentheses by 2nd parentheses